# TimeLens: Rethinking Video Temporal Grounding with Multimodal LLMs

### v1

@@ 6 Conclusion
[…]
Supplementary Material
Appendix A Details of TimeLens-Bench
@@
We construct our TimeLens-Bench based on Charades-STA [charades-sta], ActivityNet Captions [activitynet-captions], and QVHighlights [qvhighlights]. For Charades-STA and ActivityNet Captions, we utilize their test splits, while for QVHighlights, we use the validation split, as its test split annotations are not publicly available and prior works [liu2025videomind, guo2025vtg-llm] also adopt the validation split. Since the test set of ActivityNet Captions is excessively large, resulting in prohibitively high evaluation cost, we uniformly partition videos based on their duration and sample an equal number of videos from each duration bin. This yields a subset with a video count comparable to Charades-STA and QVHighlights, while maintaining a balanced distribution of video durations. Although QVHighlights was originally annotated for both video temporal grounding and video highlight detection, our work focuses exclusively on the temporal grounding task.
@@
Charades-STA [charades-sta] 1334 29.5 3720 - - Daily Life
@@
ActivityNet Captions [activitynet-captions] 4885 118.1 17031 - - Activity
@@
Table 4: Statistics of the datasets in our proposed TimeLens-Bench, compared against their original versions (Charades-STA [charades-sta], ActivityNet Captions [activitynet-captions] and QVHighlights [qvhighlights]).
@@ A.1 Evaluation Metrics
GPT-4o [hurst2024gpt-4o] 60.6 44.5 23.5 41.8 51.6 27.9 11.7 34.7
GPT-5 [OpenAI2025_GPT5] 59.3 42.0 22.0 40.5 39.7 18.3 6.2 28.4
Gemini-2.0-Flash [comanici2025gemini-2-5] 66.4 53.5 27.1 46.7 55.6 29.0 9.5 35.1
Gemini-2.5-Flash [comanici2025gemini-2-5] 68.7 56.1 30.6 48.6 47.0 21.8 7.1 31.1
Gemini-2.5-Pro [comanici2025gemini-2-5] 74.1 61.1 34.0 52.8 53.9 25.5 8.8 34.6
@@
VideoChat-Flash-7B [li2024videochat_flash] 60.2 37.9 17.8 39.7 72.5* 51.4* 26.4* 45.2*
VideoChat-R1-7B [li2025videochat_r1] 51.9 30.8 11.7 33.7 - 71.7* 50.2* 60.8*
Time-R1-7B [wang2025timer1] 57.9 32.0 16.9 36.6 78.1* 60.8* 35.3* 58.1*
MiMo-VL-7B [mimo-vl] 57.9 42.6 20.5 39.6 - - - 50.0*
Qwen2.5-VL-7B [qwen2-5-vl] (Baseline) 59.7 37.8 16.6 39.3 59.4 38.2 18.1 43.6*
@@
Appendix B More Implementation Details.
@@
Gemini-2.0-Flash [comanici2025gemini-2-5] 62.9 54.0 37.7 49.3 50.4 33.2 19.9 36.5
Gemini-2.5-Flash [comanici2025gemini-2-5] 66.8 57.5 41.3 52.5 51.2 34.7 21.0 37.4
@@
VideoChat-R1-7B [li2025videochat_r1] 35.0 23.9 11.3 25.0 - 33.3* 16.7* 35.5*
Time-R1-7B [wang2025timer1] 44.8 31.0 19.0 33.1 58.1* 39.0* 21.4* 40.5*
MiMo-VL-7B [mimo-vl] 49.3 38.7 22.4 35.5 39.3 24.3 12.9 28.1
Qwen2.5-VL-7B [qwen2-5-vl] (Baseline) 44.1 31.0 16.1 31.4 34.5 20.8 11.2 26.4
@@
Gemini-2.0-Flash [comanici2025gemini-2-5] 76.2 66.4 48.3 60.8 72.1 58.1 41.9 54.9
Gemini-2.5-Flash [comanici2025gemini-2-5] 78.2 69.4 55.0 64.3 76.9 62.6 46.7 59.1
@@
MiMo-VL-7B [mimo-vl] 57.1 42.6 28.4 41.5 59.6 43.3 24.7 41.8
Qwen2.5-VL-7B [qwen2-5-vl] 41.5 27.8 15.2 31.6 29.2 19.2 11.9 27.4
@@ B.1 Preliminaries for RL
We formalize the distinction between thinking-based and thinking-free RLVR paradigms using GRPO [shao2024deepseekmath] as the reinforcement learning algorithm.
@@
The key distinction between the two paradigms lies in the response structure and reward computation. In thinking-based RLVR, following the “think-then-answer” approach [guo2025deepseek_r1], the response consists of two parts:
@@
To formalize difficulty-aware sampling [yuan2025vl_cogito, wang2025timer1, hong2025glm_4_5], we first perform offline inference with the model to be trained on the training dataset $\mathcal{D}=\{(v_{i},q_{i},S_{i}^{*})\}_{i=1}^{N}$. For each sample, we obtain the predicted segment $\hat{S}_{i}$ and compute the difficulty estimate as:
@@
We then compute sampling weights for each sample based on its difficulty. Following  [yuan2025vl_cogito, wang2025timer1], we employ Gaussian sampling to construct a training subset where samples with difficulty around a target mean $\mu$ are more likely to be selected. Let $g(d;\mu,\sigma^{2})$ denote the target Gaussian distribution:
@@
B.2 Experimental Setup
Unless otherwise specified, all experiments are conducted using Qwen2.5-VL-7B [qwen2-5-vl] as the base model. Under the Qwen2.5-VL architecture, every two consecutive video frames are merged during the patch-embedding stage of the vision encoder.
@@ B.2 Experimental Setup
In Tab. 2 and Sec. 5.1, we experiment with different timestamp encoding methods. For position-embedding methods, we directly adopt the original MRoPE implementation from Qwen2.5-VL [qwen2-5-vl]. For Visual Overlay, we render timestamps in red with a font size of $40$ pt, overlaid at the bottom-left corner of each frame. For Non-interleaved Textual Encoding, an instruction like “This video samples $N$ frames of a $T$-second video at $t_{1},t_{2},\ldots$ seconds” is prepended to the prompt.
@@
Throughout the development of this work, our experiments were conducted based on Qwen2.5-VL. More recently, the more powerful Qwen3-VL [Qwen3-VL] models have been released, so we also validated the effectiveness of our data and recipe on Qwen3-VL. We observed that directly applying RL training on Qwen3-VL fails to yield improvements, likely because Qwen3-VL has undergone large-scale multi-task RL training that includes VTG data, preventing the model from generating rollouts with sufficient diversity on VTG task during our continual RL. Therefore, we first perform a small SFT stage to, in a sense, revert the model back to the “base model” state before RL. This is merely a workaround specific to Qwen3-VL, a model that has already acquired strong VTG capabilities through an RL stage similar to that proposed in this paper. In the common scenario, our recipes are designed to enhance the VTG capabilities of a “base MLLM”, where this trick is not required.
Appendix C More Experimental Results
@@
In Tab. 10, we evaluate TimeLens-7B’s general video understanding capabilities on VideoMME [fu2025videomme], the most comprehensive and widely-adopted video understanding benchmark. The results demonstrate that TimeLens-7B maintains the strong general video understanding capability of its base model. This validates that our proposed design principles can effectively enhance video temporal grounding capabilities without sacrificing general-purpose video understanding abilities.
@@
Table 10: Results on the general video understanding benchmark Video-MME [fu2025videomme]. The results show that TimeLens-Qwen2.5-VL-7B maintains strong general video understanding capability while achieving substantial improvements in video temporal grounding. * Results reproduced by us. Results reported in original paper.
Appendix D Discussion on Thinking-free vs. Thinking-based RLVR
@@
Appendix E Annotation Interface and Manual
@@
Appendix F Details of Curating TimeLens-100K
As described in Appendix F, we perform automated re-annotation on existing training datasets, resulting in TimeLens-100K, a large-scale, high-quality VTG training set comprising approximately 20K videos and 100K VTG annotations.
We begin by sampling videos from numerous existing VTG datasets, including CosMo-Cap [wang2024cosmo_cap], InternVid-VTime [huang2024vtimellm-internvid-vtime], DiDeMo [didemo], QuerYD [oncescu2021queryd], HiREST [zala2023hirest], etc. These datasets already cover sufficiently diverse video domains. Additionally, we perform uniform sampling based on video duration to ensure sampled videos are approximately uniformly distributed within 0–240 seconds, with a small portion of longer videos included.
@@ Appendix F Details of Curating TimeLens-100K
Specifically, we use Gemini-2.5-Pro [comanici2025gemini-2-5], currently the best-performing VTG model, for re-annotation. The annotation prompt is provided in Fig. 10. Notably, although this prompt appears simple, it is the result of extensive prompt engineering and optimization. We find that a concise and intuitive prompt is sufficient, as the model possesses adequate common sense and reasoning capabilities to understand the task. Overly complex and detailed prompts are unnecessary and can actually degrade annotation quality. During its reasoning process, the model can automatically verify and ensure the uniqueness and uniform distribution of events throughout the video.
@@
Appendix G Implementation Details for Benchmarking Existing MLLMs
@@
The prompt for training and evaluating TimeLens models is illustrated in Fig. 9. Implementation details are provided in Appendix B.
GPT-5 [OpenAI2025_GPT5] and GPT-4o [hurst2024gpt-4o]
Since GPT models only support multi-image sequences as input, we sample frames from videos at 1 FPS and prepend textual timestamps (i.e., “Frame at 2.5s:”) to each frame. As the Azure OpenAI API we use does not support more than 50 images for a single request, we adopt different strategies for videos longer than 50 seconds: for videos lasting 50-80 seconds, we uniformly sample 50 frames; for videos longer than 80 seconds, we sample at 1 FPS and arrange every 4 consecutive frames into a $2\times 2$ grid within a single image, following previous works [ju2024miradata, zhang2024internlm-xcomposer-2-5, zhang2024longva]. For GPT-5, which is a thinking model, we use the default value for the reasoning.effort parameter. The evaluation prompt is shown in Fig. 14.
Gemini models [comanici2025gemini-2-5]
@@
Qwen3-VL [Qwen3-VL], Qwen2.5-VL [qwen2-5-vl] and MiMo-VL [mimo-vl]
TimeLens models, Qwen2.5-VL-7B, and MiMo-VL share approximately the same model architecture and hyperparameter configurations. Therefore, when evaluating these models, we adopt the same settings to ensure fair comparisons in Tab. 1. Specifically, consistent with Sec. B.2, we sample video frames at 2 FPS and set the resolution budget to $\texttt{min\_tokens}=64$ and $\texttt{total\_tokens}=14,336$. For MiMo-VL, we evaluate their best-performing model, MiMo-VL-7B-RL. The evaluation prompt is shown in Fig. 14.
@@
When evaluating Time-R1 [wang2025timer1], VideoChat-Flash [li2024videochat_flash] and VideoChat-R1 [li2025videochat_r1], we directly use their original codebases. Please refer to their papers and code repositories for details.

### v2

@@ Appendix A Conclusion
Appendix B Details of TimeLens-Bench
@@
We construct our TimeLens-Bench based on Charades-STA [11], ActivityNet Captions [25], and QVHighlights [27]. For Charades-STA and ActivityNet Captions, we utilize their test splits, while for QVHighlights, we use the validation split, as its test split annotations are not publicly available and prior works [36, 17] also adopt the validation split. Since the test set of ActivityNet Captions is excessively large, resulting in prohibitively high evaluation cost, we uniformly partition videos based on their duration and sample an equal number of videos from each duration bin. This yields a subset with a video count comparable to Charades-STA and QVHighlights, while maintaining a balanced distribution of video durations. Although QVHighlights was originally annotated for both video temporal grounding and video highlight detection, our work focuses exclusively on the temporal grounding task.
@@
Charades-STA [11] 1334 29.5 3720 - - Daily Life
@@
ActivityNet Captions [25] 4885 118.1 17031 - - Activity
@@
Table 4: Statistics of the datasets in our proposed TimeLens-Bench, compared against their original versions (Charades-STA [11], ActivityNet Captions [25] and QVHighlights [27]).
@@ B.1 Evaluation Metrics
GPT-4o [23] 60.6 44.5 23.5 41.8 51.6 27.9 11.7 34.7
GPT-5 [42] 59.3 42.0 22.0 40.5 39.7 18.3 6.2 28.4
Gemini-2.0-Flash [8] 66.4 53.5 27.1 46.7 55.6 29.0 9.5 35.1
Gemini-2.5-Flash [8] 68.7 56.1 30.6 48.6 47.0 21.8 7.1 31.1
Gemini-2.5-Pro [8] 74.1 61.1 34.0 52.8 53.9 25.5 8.8 34.6
@@
VideoChat-Flash-7B [31] 60.2 37.9 17.8 39.7 72.5* 51.4* 26.4* 45.2*
VideoChat-R1-7B [32] 51.9 30.8 11.7 33.7 - 71.7* 50.2* 60.8*
Time-R1-7B [53] 57.9 32.0 16.9 36.6 78.1* 60.8* 35.3* 58.1*
MiMo-VL-7B [9] 57.9 42.6 20.5 39.6 - - - 50.0*
Qwen2.5-VL-7B [3] (Baseline) 59.7 37.8 16.6 39.3 59.4 38.2 18.1 43.6*
@@
Appendix C More Implementation Details.
@@
Gemini-2.0-Flash [8] 62.9 54.0 37.7 49.3 50.4 33.2 19.9 36.5
Gemini-2.5-Flash [8] 66.8 57.5 41.3 52.5 51.2 34.7 21.0 37.4
@@
VideoChat-R1-7B [32] 35.0 23.9 11.3 25.0 - 33.3* 16.7* 35.5*
Time-R1-7B [53] 44.8 31.0 19.0 33.1 58.1* 39.0* 21.4* 40.5*
MiMo-VL-7B [9] 49.3 38.7 22.4 35.5 39.3 24.3 12.9 28.1
Qwen2.5-VL-7B [3] (Baseline) 44.1 31.0 16.1 31.4 34.5 20.8 11.2 26.4
@@
Gemini-2.0-Flash [8] 76.2 66.4 48.3 60.8 72.1 58.1 41.9 54.9
Gemini-2.5-Flash [8] 78.2 69.4 55.0 64.3 76.9 62.6 46.7 59.1
@@
MiMo-VL-7B [9] 57.1 42.6 28.4 41.5 59.6 43.3 24.7 41.8
Qwen2.5-VL-7B [3] 41.5 27.8 15.2 31.6 29.2 19.2 11.9 27.4
@@ C.1 Preliminaries for RL
We formalize the distinction between thinking-based and thinking-free RLVR paradigms using GRPO [47] as the reinforcement learning algorithm.
@@
The key distinction between the two paradigms lies in the response structure and reward computation. In thinking-based RLVR, following the “think-then-answer” approach [15], the response consists of two parts:
@@
To formalize difficulty-aware sampling [57, 53, 21], we first perform offline inference with the model to be trained on the training dataset $\mathcal{D}=\{(v_{i},q_{i},S_{i}^{*})\}_{i=1}^{N}$. For each sample, we obtain the predicted segment $\hat{S}_{i}$ and compute the difficulty estimate as:
@@
We then compute sampling weights for each sample based on its difficulty. Following  [57, 53], we employ Gaussian sampling to construct a training subset where samples with difficulty around a target mean $\mu$ are more likely to be selected. Let $g(d;\mu,\sigma^{2})$ denote the target Gaussian distribution:
@@
C.2 Experimental Setup
Unless otherwise specified, all experiments are conducted using Qwen2.5-VL-7B [3] as the base model. Under the Qwen2.5-VL architecture, every two consecutive video frames are merged during the patch-embedding stage of the vision encoder.
@@ C.2 Experimental Setup
In Tab. 2 and Sec. 5.1, we experiment with different timestamp encoding methods. For position-embedding methods, we directly adopt the original MRoPE implementation from Qwen2.5-VL [3]. For Visual Overlay, we render timestamps in red with a font size of $40$ pt, overlaid at the bottom-left corner of each frame. For Non-interleaved Textual Encoding, an instruction like “This video samples $N$ frames of a $T$-second video at $t_{1},t_{2},\ldots$ seconds” is prepended to the prompt.
@@
Throughout the development of this work, our experiments were conducted based on Qwen2.5-VL. More recently, the more powerful Qwen3-VL [2] models have been released, so we also validated the effectiveness of our data and recipe on Qwen3-VL. We observed that directly applying RL training on Qwen3-VL fails to yield improvements, likely because Qwen3-VL has undergone large-scale multi-task RL training that includes VTG data, preventing the model from generating rollouts with sufficient diversity on VTG task during our continual RL. Therefore, we first perform a small SFT stage to, in a sense, revert the model back to the “base model” state before RL. This is merely a workaround specific to Qwen3-VL, a model that has already acquired strong VTG capabilities through an RL stage similar to that proposed in this paper. In the common scenario, our recipes are designed to enhance the VTG capabilities of a “base MLLM”, where this trick is not required.
Appendix D More Experimental Results
@@
Generalization to an external benchmark
To further validate that our recipe generalizes beyond TimeLens-Bench, we additionally evaluate on VUE-TR, a high-quality human-annotated VTG benchmark introduced concurrently in Vidi [49]. As shown in Tab. 10, TimeLens-7B achieves the best result among the compared models, indicating that the gains from our data curation and algorithmic recipe are not limited to our own refined benchmark suite.
Model IoU @ VUE-TR (0-200s)
Qwen2.5-VL-7B [3] 36.0
GPT-4o [23] 34.5
Gemini-2.5-Pro [8] 41.6
[…]
Table 10: Results on the VUE-TR benchmark from Vidi [49]. TimeLens-7B achieves the best IoU among the compared models on this external benchmark.
@@
In Tab. 11, we evaluate TimeLens-7B’s general video understanding capabilities on VideoMME [10], the most comprehensive and widely-adopted video understanding benchmark. The results demonstrate that TimeLens-7B maintains the strong general video understanding capability of its base model. This validates that our proposed design principles can effectively enhance video temporal grounding capabilities without sacrificing general-purpose video understanding abilities.
@@
Table 11: Results on the general video understanding benchmark Video-MME [10]. The results show that TimeLens-Qwen2.5-VL-7B maintains strong general video understanding capability while achieving substantial improvements in video temporal grounding. * Results reproduced by us. Results reported in original paper.
Appendix E Discussion on Thinking-free vs. Thinking-based RLVR
@@
Appendix F Additional Discussions
Why do VideoChat-R1 and Time-R1 behave differently across benchmarks?
VideoChat-R1 [32] is trained on single-domain data with relatively limited quality control, and it does not incorporate the training recipes that we find most effective for VTG, such as difficulty-aware sampling and early stopping. As a result, it underperforms even the Qwen2.5-VL baseline on our refined benchmarks. Time-R1 [53] by contrast, benefits from a stronger RL-oriented training recipe and therefore transfers better to ActivityNet-TimeLens and QVHighlights-TimeLens, where long-video localization is a major challenge. However, its training data is still noisier than ours, which likely limits temporal precision and makes it less competitive on Charades-TimeLens, where videos are shorter but localization accuracy must be much higher.
Training setting for timestamp encoding.
All methods in Tab. 2 are compared after RLVR training rather than in a training-free setting. This controlled setup is important because the temporal encoding method interacts with the optimization paradigm: the final comparison reflects how well each representation supports post-training for VTG, instead of only measuring zero-shot prompting behavior.
Early stopping.
In our single-task VTG setting, reward plateauing is a reliable signal that continued RLVR is unlikely to improve grounding quality and may even degrade it, so early stopping saves computation while preventing over-training. We regard this as a practical recipe rather than a universal rule. In more general multi-task training, a single stopping criterion may be harder to define because different capabilities can peak at different stages, requiring more sophisticated recipes (e.g., remove training data for a single task when reward on this task peaks, while continuing training for other tasks).
Appendix G Annotation Interface and Manual
@@
Appendix H Details of Curating TimeLens-100K
As described in Appendix H, we perform automated re-annotation on existing training datasets, resulting in TimeLens-100K, a large-scale, high-quality VTG training set comprising approximately 20K videos and 100K VTG annotations.
We begin by sampling videos from numerous existing VTG datasets, including CosMo-Cap [50], InternVid-VTime [22], DiDeMo [1], QuerYD [41], HiREST [60], etc. These datasets already cover sufficiently diverse video domains. Additionally, we perform uniform sampling based on video duration to ensure sampled videos are approximately uniformly distributed within 0–240 seconds, with a small portion of longer videos included.
@@ Appendix H Details of Curating TimeLens-100K
Specifically, we use Gemini-2.5-Pro [8], currently the best-performing VTG model, for re-annotation. The annotation prompt is provided in Fig. 10. Notably, although this prompt appears simple, it is the result of extensive prompt engineering and optimization. We find that a concise and intuitive prompt is sufficient, as the model possesses adequate common sense and reasoning capabilities to understand the task. Overly complex and detailed prompts are unnecessary and can actually degrade annotation quality. During its reasoning process, the model can automatically verify and ensure the uniqueness and uniform distribution of events throughout the video.
@@
Appendix I Implementation Details for Benchmarking Existing MLLMs
@@
The prompt for training and evaluating TimeLens models is illustrated in Fig. 9. Implementation details are provided in Appendix C.
GPT-5 [42] and GPT-4o [23]
Since GPT models only support multi-image sequences as input, we sample frames from videos at 1 FPS and prepend textual timestamps (i.e., “Frame at 2.5s:”) to each frame. As the Azure OpenAI API we use does not support more than 50 images for a single request, we adopt different strategies for videos longer than 50 seconds: for videos lasting 50-80 seconds, we uniformly sample 50 frames; for videos longer than 80 seconds, we sample at 1 FPS and arrange every 4 consecutive frames into a $2\times 2$ grid within a single image, following previous works [24, 63, 64]. For GPT-5, which is a thinking model, we use the default value for the reasoning.effort parameter. The evaluation prompt is shown in Fig. 14.
Gemini models [8]
@@
Qwen3-VL [2], Qwen2.5-VL [3] and MiMo-VL [9]
TimeLens models, Qwen2.5-VL-7B, and MiMo-VL share approximately the same model architecture and hyperparameter configurations. Therefore, when evaluating these models, we adopt the same settings to ensure fair comparisons in Tab. 1. Specifically, consistent with Sec. C.2, we sample video frames at 2 FPS and set the resolution budget to $\texttt{min\_tokens}=64$ and $\texttt{total\_tokens}=14,336$. For MiMo-VL, we evaluate their best-performing model, MiMo-VL-7B-RL. The evaluation prompt is shown in Fig. 14.
@@
When evaluating Time-R1 [53], VideoChat-Flash [31] and VideoChat-R1 [32], we directly use their original codebases. Please refer to their papers and code repositories for details.